# High-Level Multi-Robot Trajectory Planning And Spurious Behavior Detection

### v1

@@ III-A Formal specification of multi-robot missions via NWNs
Finally, top right of the figure shows the System Net. This high-level Petri net contains tokens that represent the previously described component nets: three tokens in place $Rb$, each of these tokens correspond to a robot net and another token in place $Ms$ corresponding to the mission specification net. Also for each of these transitions, a different number of robots can move synchronously, e.g., one, two or three robots moving at the same time in the example net. Transitions in this system net are responsible for synchronizing the evolution of the robot models with the mission specification. The synchronization is managed via the Global Enabling Function (GEF), Java functions integrated into the system net (see [13] for further details).
@@ III-B Dataset generation: simulating trajectories with Renew
The algorithm for trajectory generation is formally initialized with all the above-defined Petri nets where initial markings are configured based on the selected environment. An empty buffer is then instantiated to record the generated trajectory steps. The system then enters an iterative loop that continues until a designated final place, $p_{end}$, in the Specifiaction Net ($\mathcal{N}^{S}$) is marked. Within each loop iteration, the system evaluates all transitions currently enabled in $\mathcal{N}^{S}$. For each of these, the Global Enabling Function ($GEF$) identifies all possible combinations of robot transitions that are enabled in the respective Robot nets ($\mathcal{N}^{o^{j}}$). A valid macro-transition is then selected, utilizing a random policy to ensure the generation of rich and varied trajectories. Firing this transition atomically updates the markings in both the $\mathcal{N}^{S}$ and the respective $\mathcal{N}^{o^{j}}$ nets. These transitions are subsequently recorded in the trajectory buffer for the current step. The execution terminates either when the $GEF$ cannot identify any possible robot transitions for the current state, signifying that the mission is impossible to complete within the given environment (e.g., the goal is completely blocked by obstacles) or when, as explained, $p_{end}$ has been reached.
@@
These symbolic logs form the raw input for the preprocessing stage, where the trajectories are converted into numerical tensors for the downstream machine-learning pipeline. The Global Enabling Function ($GEF$) and robot-specific subnets facilitate the controlled injection of spurious behaviors, including coordination failures, timing mismatches, or task divergence. Following a similar generation process used for correct trajectories, spurious datasets are created by modifying either the robot nets to alter scenarios or introduce timing mismatches, or by subtly adjusting the specification net to induce subtle high-level behavioral deviations from the intended mission.

### v2

@@ III-A Formal specification of multi-robot missions via NWNs
Finally, top right of the figure shows the System Net. This high-level Petri net contains tokens that represent the previously described component nets: three tokens in place RbRb, each of these tokens correspond to a robot net and another token in place MsMs corresponding to the mission specification net. Also for each of these transitions, a different number of robots can move synchronously, e.g., one, two or three robots moving at the same time in the example net. Transitions in this system net are responsible for synchronizing the evolution of the robot models with the mission specification. The synchronization is managed via the Global Enabling Function (GEF), Java functions integrated into the system net (see [13] for further details).
@@ III-B Dataset generation: simulating trajectories with Renew
The algorithm for trajectory generation is formally initialized with all the above-defined Petri nets where initial markings are configured based on the selected environment. An empty buffer is then instantiated to record the generated trajectory steps. The system then enters an iterative loop that continues until a designated final place, $p_{end}$, in the Specifiaction Net ($\mathcal{N}^{S}$) is marked. Within each loop iteration, the system evaluates all transitions currently enabled in $\mathcal{N}^{S}$. For each of these, the Global Enabling Function (GEFGEF) identifies all possible combinations of robot transitions that are enabled in the respective Robot nets ($\mathcal{N}^{o^{j}}$). A valid macro-transition is then selected, utilizing a random policy to ensure the generation of rich and varied trajectories. Firing this transition atomically updates the markings in both the $\mathcal{N}^{S}$ and the respective $\mathcal{N}^{o^{j}}$ nets. These transitions are subsequently recorded in the trajectory buffer for the current step. The execution terminates either when the GEFGEF cannot identify any possible robot transitions for the current state, signifying that the mission is impossible to complete within the given environment (e.g., the goal is completely blocked by obstacles) or when, as explained, $p_{end}$ has been reached.
@@
These symbolic logs form the raw input for the preprocessing stage, where the trajectories are converted into numerical tensors for the downstream machine-learning pipeline. The Global Enabling Function (GEFGEF) and robot-specific subnets facilitate the controlled injection of spurious behaviors, including coordination failures, timing mismatches, or task divergence. Following a similar generation process used for correct trajectories, spurious datasets are created by modifying either the robot nets to alter scenarios or introduce timing mismatches, or by subtly adjusting the specification net to induce subtle high-level behavioral deviations from the intended mission.